# StrokeFusion: Vector Sketch Generation via Joint Stroke-UDF Encoding and Latent Sequence Diffusion

### v3

@@ Abstract
In the field of sketch generation, raster-format trained models often produce non-stroke artifacts, while vector-format trained models typically lack a holistic understanding of sketches, leading to compromised recognizability. Moreover, existing methods struggle to extract common features from similar elements (e.g., eyes of animals) appearing at varying positions across sketches. To address these challenges, we propose StrokeFusion, a two-stage framework for vector sketch generation. It contains a dual-modal sketch feature learning network that maps strokes into a high-quality latent space. This network decomposes sketches into normalized strokes and jointly encodes stroke sequences with Unsigned Distance Function (UDF) maps, representing sketches as sets of stroke feature vectors. Building upon this representation, our framework exploits a stroke-level latent diffusion model that simultaneously adjusts stroke position, scale, and trajectory during generation. This enables high-fidelity stroke generation while supporting stroke interpolation editing. Extensive experiments across multiple sketch datasets demonstrate that our framework outperforms state-of-the-art techniques, validating its effectiveness in preserving structural integrity and semantic features. Code and models will be made publicly available upon publication.
@@ Introduction
Current sketch representations primarily exist in two formats: raster sketches and vector sketches. Raster sketches render strokes directly as images and are learned through reconstruction of raster images (wang2024vq; ge2020creative; bhunia2022doodleformer), allowing the use of mature image representation learning methods such as CNN and VQ-VAE (van2017neural). However, these models often produce pixelation artifacts and non-stroke noise. More critically, they discard stroke trajectories, making it difficult to capture human drawing habits and natural stroke patterns, focusing solely on overall shape.
In contrast, vector sketches preserve clean and precise stroke trajectories that better simulate human drawing processes. Current vector-based approaches typically trace the entire trajectory of strokes as poly-lines, recording pen-up and pen-down states. This leads to two modeling paradigms: modeling strokes using velocity-based relative coordinates (ha2017neural; das2023chirodiff), or using absolute position representations (wang2023sketchknitter; bandyopadhyay2024sketchinr). Relative coordinate modeling allows the model to learn common patterns more easily (e.g., eyes, wheels), since velocity vectors are often similar across these structures. However, these models must integrate velocity vectors to determine current positions, causing error accumulation that makes it difficult to form closed loops or position components accurately. Absolute coordinate modeling alleviates these issues but weakens the model’s ability to capture stroke-wise commonalities, leading to inferior quality compared to velocity-based methods.
@@
Unlike vector-based approaches, several methods aim to generate sketches directly in raster form. DoodlerGAN (ge2020creative) introduced a part-based generative adversarial framework that learns the appearance and composition of semantic parts (e.g., head, wing) purely in image space. DoodleFormer (bhunia2022doodleformer) further enhanced structural understanding by proposing a coarse-to-fine generation strategy using a two-stage Transformer architecture. VQ-SGen  (wang2024vq) employed vector quantization to encode sketches into discrete stroke patches, which are then decoded using an autoregressive Transformer. These methods achieve visually compelling results and demonstrate strong structure modeling compared to point-based models. However, they lack the ability to produce editable vector outputs. In contrast, our method leverages the shape modeling strengths of image-level supervision while retaining the benefits of vector-format outputs.
@@
Ha et al. (ha2017neural) introduced stroke-annotated polyline sequences along with the foundational Sketch-RNN model. Ribeiro et al. (ribeiro2020sketchformer), Xu et al. (xu2024sketchformer++), and Lin et al. (lin2020sketch) extended Transformer-based architectures by integrating local-global attention and self-supervised BERT-like objectives for recognition and completion. While effective for representation learning, these works primarily target downstream tasks rather than generative modeling.
Qi et al. (qi2022generative) adopted Graph Convolutional Networks (GCNs) (kipf2016semi) to model local structures through sketch partitioning. Tiwari et al. (tiwari2024sketchgpt) tokenized sketches into stroke dictionaries for autoregressive GPT-style decoding. Zang et al. (zang2023self) employed Gaussian mixture-based latent spaces to model structure-aware sketch priors. Wang et al. (wang2023sketchknitter) modeled vector sketches as offset sequences with intermediate recognizability scoring to improve denoising diffusion efficiency. Das et al. (das2023chirodiff) framed velocity fields as continuous-time functions and employed non-autoregressive diffusion for temporally coherent trajectory synthesis. However, these methods treat sketches as point sequences interleaved with pen states, which encourages models to overfit to superficial variations like stroke density or sampling intervals. This leads to inefficient training and poor generalization across different drawing styles. Inspired by recent progress in 3D shape generation (zhang20233dshape2vecset), our method encodes each stroke independently into fixed-length feature vectors and treats them as an unordered variable-length set during training, accommodating the intrinsic variability in stroke count and order.
@@
Traditional methods often discretize sketches into points or linear segments (ha2017neural), ignoring the fact that human drawing is inherently continuous in both time and space. To bridge this gap, several approaches explore continuous representations to better capture local deformations and global semantics. PMN (alaniz2022abstracting) represented strokes using primitives from a fixed graphical library. BézierSketch (das2020beziersketch) and Cloud2Curve (das2021cloud2curve) used Bézier curves, with the latter introducing adaptive order to increase flexibility. Stroke Clouds (ashcroft2023modelling) segmented trajectories into third-order Bézier curves and used diffusion models to learn their distribution. Nonetheless, these techniques still inherit the inductive bias and expressiveness bottleneck of Bézier primitives.
Recent efforts have shifted toward learning fully continuous functions. SketchODE (das2022sketchode) modeled stroke dynamics using Neural Ordinary Differential Equations (ODEs), learning a time-evolving latent path. SketchINR (bandyopadhyay2024sketchinr) represented sketches as continuous neural fields (Implicit Neural Representations), mapping time to coordinates. While promising, these methods frequently overemphasize easily fitted components, neglecting holistic coherence. Our approach encodes entire strokes as learnable embeddings with associated stroke distance fields, which allow the generative model to reason about their shapes during layout synthesis. This enables more semantically coherent sketch composition while preserving the ability to recover clean vector trajectories.
@@
Our work builds on Diffusion Models (sohl2015deep), which operate through forward (gradual noising) and reverse (iterative denoising) processes. Recent advances include continuous-space implementations (dhariwal2021diffusion; ho2020denoising) and discrete-space extensions like D3PM (austin2021structured). Our method builds on these advancements in diffusion modeling while drawing inspiration from recent progress in 3D shape representation (zhang20233dshape2vecset; zhang2024clay). We adapt this vector-set encoding approach to represent strokes, enabling diffusion-based generation of drawing primitives through iterative denoising processes. This combination of discrete-state diffusion techniques with structured vector representations allows our model to benefit from both robust probabilistic modeling and geometrically meaningful feature learning.
@@
Figure 2 illustrates two representative examples of the denoising process. As the noise level decreases, each stroke progressively refines its shape, spatial placement, and presence confidence by implicitly referencing other strokes within the composition. This self-organizing behavior emerges from the model’s joint reasoning over the entire stroke set.
Unlike many existing vector sketch generation methods that treat all strokes as a single concatenated trajectory (yang2021sketchgnn; das2023chirodiff; wang2023sketchknitter), our design allows for a flexible number of strokes and diverse layout patterns. This is achieved by processing strokes as an unordered set without relying on positional encodings, thereby preserving permutation invariance throughout the generation process.
@@
SketchRNN 31.607 0.490 0.449 36.978 0.580 0.439 40.667 0.552 0.396
SketchKnitter 23.166 0.518 0.479 27.067 0.571 0.450 35.635 0.541 0.401
ChiroDiff 17.166 0.607 0.496 23.838 0.630 0.452 27.784 0.618 0.418
StrokeFusion (Ours) 19.532 0.708 0.581 18.992 0.689 0.607 17.762 0.710 0.583
@@
We compare our method with three representative sketch generation models: SketchRNN (yang2021sketchgnn), SketchKnitter (wang2023sketchknitter), and ChiroDiff (das2023chirodiff). These methods all focus on vector sketch generation rather than traditional raster image modeling and can directly output stroke sequences.
SketchRNN is an RNN-based variational autoencoder that can generate stroke sequences under conditional or unconditional settings and control generation diversity via a temperature parameter. SketchKnitter introduces a diffusion model framework, treating the generation of stroke points as a progressive denoising process from random initial noise to a clear sketch, significantly improving structural fidelity. ChiroDiff further incorporates a temporal modeling mechanism by resampling strokes, thereby better capturing stroke order and dynamic changes during generation.
@@
FID measures the distributional distance between real and generated data in the Inception feature space, with lower scores indicating better visual fidelity. Precision reflects the proportion of generated samples close to the real data manifold (fidelity), while Recall captures how well the generated samples cover the diversity of real data. These complementary metrics jointly assess the perceptual quality and coverage of the model.
@@
SketchRNN 59.859 0.260 0.278 121.020 0.438 0.256 155.020 0.010 0.312 98.006 0.730 0.202
SketchKnitter 59.159 0.252 0.216 110.456 0.416 0.267 156.972 0.075 0.335 99.462 0.564 0.221
ChiroDiff 60.104 0.561 0.175 36.656 0.588 0.272 99.326 0.062 0.303 98.301 0.528 0.253
Doodleformer 27.322 0.671 0.549 33.456 0.523 0.685 - - - - - -
StrokeFusion (Ours) 26.191 0.557 0.303 19.407 0.579 0.315 7.265 0.755 0.891 33.683 0.657 0.480
@@
To better showcase structural modeling ability of StrokeFusion, we also evaluate on four more challenging datasets: Creative Birds, Creative Creatures, FaceX, and TU Berlin. As shown in Table 2, our method consistently outperforms all baselines across metrics, confirming strong generalization across styles and domains.
@@
w/o stroke normalization 46.733 0.669 0.035 105.648 0.242 0.030 84.443 0.614 0.021
w/o image-level supervision 38.027 0.568 0.268 46.519 0.532 0.134 139.279 0.112 0.001
The full model 24.998 0.678 0.414 16.276 0.630 0.549 15.185 0.708 0.480
@@
For Creative Birds and Creative Creatures, the high stylistic variation and imaginative designs pose significant challenges. Diffusion-based baselines capture coarse contours but struggle with detail, while SketchRNN suffers from degraded quality in longer sequences. Our method balances structure and detail more effectively, yielding clear and recognizable outputs.
@@ Conclusion
We have introduced StrokeFusion, a novel framework for vector sketch generation that integrates dual-modal stroke encoding with continuous diffusion modeling. Our dual-branch encoder jointly captures geometric and visual features from point sequences and stroke distance fields, fusing them into a unified latent representation. This enriched embedding space facilitates smooth feature transitions during diffusion, enabling the generation of sketches with natural stroke fluidity, structural coherence, and fine-grained detail. Moreover, our method preserves spatial relationships throughout the denoising process and efficiently reconstructs realistic drawing trajectories. Extensive qualitative and quantitative evaluations demonstrate the superiority of StrokeFusionin producing diverse, high-quality sketches across varying levels of complexity.
@@
Figures 9 and 10 present additional qualitative results on both QuickDraw and several more complex sketch datasets. These visualizations further demonstrate the robustness and generalizability of our model across diverse domains.

### v4

@@ Abstract
In the field of sketch generation, raster-format-trained models often produce non-stroke artifacts, while vector-format-trained models typically lack a holistic understanding of sketches, resulting in compromised recognizability. Moreover, existing methods struggle to extract common features from similar elements (e.g., animal eyes) that appear at varying positions across sketches. To address these challenges, we propose StrokeFusion, a two-stage framework for vector sketch generation. It contains a dual-modal sketch feature learning network that maps strokes into a high-quality latent space. This network decomposes sketches into normalized strokes and jointly encodes stroke sequences with Unsigned Distance Function (UDF) maps, representing sketches as sets of stroke feature vectors. Building upon this representation, our framework exploits a stroke-level latent diffusion model that simultaneously adjusts stroke position, scale, and trajectory during generation. This enables high-fidelity stroke generation while supporting stroke interpolation editing. Extensive experiments across multiple sketch datasets demonstrate that our framework outperforms state-of-the-art techniques, validating its effectiveness in preserving structural integrity and semantic features. Code and models will be made publicly available upon publication.
@@ Introduction
Current sketch representations primarily exist in two formats: raster sketches and vector sketches. Raster sketches render strokes directly as images and are learned through reconstruction of raster images (wang2024vq; ge2020creative; bhunia2022doodleformer), allowing the use of mature image representation learning methods. However, these models often produce pixelation artifacts and non-stroke noise. More critically, they discard stroke trajectories, making it difficult to capture human drawing habits and natural stroke patterns, focusing solely on overall shape.
In contrast, vector sketches preserve clean and precise stroke trajectories that better simulate human drawing processes. Current vector-based approaches typically trace the entire stroke trajectory as polylines, recording pen-up and pen-down states. This leads to two modeling paradigms: modeling strokes using velocity-based relative coordinates (ha2017neural; das2023chirodiff), or using absolute position representations (wang2023sketchknitter; bandyopadhyay2024sketchinr). Relative coordinate modeling allows the model to learn common patterns more easily (e.g., eyes, wheels), since velocity vectors are often similar across these structures. However, these models must integrate velocity vectors to determine current positions, leading to error accumulation that makes it difficult to form closed loops or to determine position components accurately. Absolute coordinate modeling alleviates these issues but weakens the model’s ability to capture stroke-wise commonalities, leading to inferior quality compared to velocity-based methods.
@@
Various methods focus on generating raster sketches. DoodlerGAN (ge2020creative) introduced a part-based GAN framework for learning the appearance of semantic parts in image space. DoodleFormer (bhunia2022doodleformer) improved structural understanding using a coarse-to-fine generation strategy with a two-stage Transformer. VQ-SGen (wang2024vq) implemented vector quantization to encode sketches into discrete stroke patches for autoregressive decoding. While these methods yield visually appealing results, they do not produce editable vector outputs. Our approach combines image-level supervision in shape modeling with the benefits of vector-format outputs.
@@
Ha et al. (ha2017neural) developed stroke-annotated polyline sequences with the Sketch-RNN model. Ribeiro et al. (ribeiro2020sketchformer), Xu et al. (xu2024sketchformer++), and Lin et al. (lin2020sketch) enhanced Transformer-based architectures for representation learning aimed at recognition and completion rather than generative modeling. Qi et al. (qi2022generative) used Graph Convolutional Networks (kipf2016semi) for local structure modeling. Tiwari et al. (tiwari2024sketchgpt) tokenized sketches for GPT-style decoding, while Zang et al. (zang2023self) used Gaussian mixture latent spaces for structure-aware priors. Wang et al. (wang2023sketchknitter) represented vector sketches as offset sequences to improve diffusion efficiency. Das et al. (das2023chirodiff) approached stroke generation as continuous functions. However, many approaches overfit to superficial details, leading to inefficient training and poor generalization. Inspired by recent progress in 3D shape generation (zhang20233dshape2vecset), our method encodes strokes into fixed-length feature vectors, treating them as unordered sets to handle variability in stroke count and order.
@@
Traditional methods often discretize sketches, overlooking the continuous nature of human drawing. To address this, several methods explore continuous representations. PMN (alaniz2022abstracting) used primitives from a graphical library, while BézierSketch (das2020beziersketch) and Cloud2Curve (das2021cloud2curve) employed Bézier curves with adaptive orders. Stroke Clouds (ashcroft2023modelling) used diffusion models for third-order Bézier curves, but these techniques suffer from limited expressiveness.
Recent work focuses on fully continuous representations. SketchODE (das2022sketchode) applied Neural ODEs for stroke dynamics, and SketchINR (bandyopadhyay2024sketchinr) used implicit neural representations. However, these methods can miss holistic coherence. Our approach encodes entire strokes as learnable embeddings, enabling semantically coherent sketch composition while maintaining clean vector trajectories.
@@
Our work is rooted in Diffusion Models (sohl2015deep), which involve forward and reverse processes. Recent advances include continuous-space implementations (dhariwal2021diffusion; ho2020denoising) and discrete-space extensions like D3PM (austin2021structured). We incorporate these advancements in diffusion modeling and align them with recent progress in 3D shape representation.
@@
Figure 2 illustrates two examples of the denoising process. As the noise level decreases, each stroke progressively refines its shape, spatial placement, and presence confidence by implicitly referencing other strokes within the composition. This self-organizing behavior emerges from the model’s joint reasoning over the entire stroke set.
Unlike existing vector sketch generation methods that treat all strokes as a concatenated trajectory (yang2021sketchgnn; das2023chirodiff; wang2023sketchknitter), our design allows for a flexible number of strokes and diverse layout patterns. This is achieved by processing strokes as an unordered set without relying on positional encodings, thereby preserving permutation invariance throughout the generation process.
@@
SketchRNN 31.61 0.49 0.45 36.98 0.58 0.44 40.67 0.55 0.40
SketchKnitter 23.17 0.52 0.48 27.07 0.57 0.45 35.64 0.54 0.40
ChiroDiff 17.17 0.61 0.50 23.84 0.63 0.45 27.78 0.62 0.42
StrokeFusion 19.53 0.71 0.58 18.99 0.69 0.61 17.76 0.71 0.58
@@
We compare our method with three representative vector sketch generation models: SketchRNN (yang2021sketchgnn), SketchKnitter (wang2023sketchknitter), and ChiroDiff (das2023chirodiff). These three methods all focus on vector sketch generation, avoiding the limitations of traditional raster image modeling by directly outputting stroke sequences. Additionally, we include the set-based raster method, Doodleformer (bhunia2022doodleformer), for comparison.
SketchRNN is an RNN-based variational autoencoder that can generate stroke sequences under conditional or unconditional settings and control generation diversity via a temperature parameter. SketchKnitter introduces a diffusion model framework, treating the generation of stroke points as a progressive denoising process from random initial noise to a clear sketch, significantly improving structural fidelity. ChiroDiff further incorporates a temporal modeling mechanism by resampling strokes, thereby better capturing stroke order and dynamic changes during generation. In contrast, Doodleformer is a raster-based model that predicts bounding boxes for semantic parts (e.g., head/body) and composes them into a final raster image. Its core limitations are that it does not output controllable stroke trajectories and is fundamentally restricted to datasets with semantic part annotations, such as the Creative datasets.
@@
FID measures the distributional distance between real and generated data in the Inception feature space, with lower scores indicating better visual fidelity. Precision reflects the proportion of generated samples that are close to the real data manifold (fidelity), while Recall captures how well the generated samples capture the diversity of the real data. These complementary metrics jointly assess the model’s perceptual quality and coverage.
@@
SketchRNN 59.85 0.26 0.28 121.02 0.44 0.26 155.02 0.01 0.31 98.01 0.73 0.20
SketchKnitter 59.16 0.25 0.22 110.46 0.42 0.27 156.97 0.08 0.34 99.46 0.56 0.22
ChiroDiff 60.10 0.56 0.18 36.66 0.59 0.27 99.33 0.06 0.30 98.30 0.53 0.25
Doodleformer 27.32 0.67 0.55 33.46 0.52 0.69 - - - - - -
StrokeFusion 26.19 0.56 0.30 19.41 0.58 0.32 7.27 0.76 0.89 33.68 0.66 0.48
@@
To better showcase the structural modeling ability of StrokeFusion, we also evaluate on four more challenging datasets: Creative Birds, Creative Creatures, FaceX, and TU Berlin. As shown in Table 2, our method outperforms all baselines in general, confirming strong generalization across styles and domains.
@@
w/o stroke normalization 46.73 0.67 0.04 105.65 0.24 0.03 84.44 0.61 0.02
w/o image-level supervision 38.03 0.57 0.27 46.52 0.53 0.13 139.28 0.11 0.001
The full model 25.00 0.68 0.41 16.28 0.63 0.55 15.19 0.71 0.48
@@
Figure 5: Comparison with DoodleFormer. For each dataset we select visually similar samples from both methods.
@@
For Creative Birds and Creative Creatures, the high stylistic variation and imaginative designs pose significant challenges. Diffusion-based baselines capture coarse contours but struggle with detail, while SketchRNN suffers from degraded quality in longer sequences. Our method balances structure and detail more effectively, yielding clear and recognizable outputs. We also visually compare our results with Doodleformer (bhunia2022doodleformer), which relies on semantic part labels to generate raster sketches (see Figure 5). Our approach achieves comparable visual quality while fundamentally preserving vector representations and stroke-level controllability, unlike the raster output of Doodleformer.
@@
Limitations
While our method demonstrates strong performance, some limitations persist. First, complex strokes with high-frequency details remain challenging due to their inherent data intensity as point sequences, leading to suboptimal learning efficiency and generation quality for low-stroke sketches. Second, the unsigned distance field (UDF) decoding process introduces computational overhead, as we ultimately rely on point sequences for stroke representation. Future work should explore efficient implicit field decoding schemes or alternative UDF representations to streamline this component.
Additionally, since our method generates sketches with strokes as the minimal unit, it requires that the strokes in the dataset contain semantic information. Vector sketches generated by algorithms often feature strokes that span multiple semantic parts. This structure prevents our model from producing high-quality results. A potential improvement would be to segment input sketches in a more flexible, symbolically semantic manner, allowing better adaptation to such cases.
@@ Conclusion
We have introduced StrokeFusion, a novel framework for vector sketch generation that integrates dual-modal stroke encoding with continuous diffusion modeling. Our dual-branch encoder jointly captures geometric and visual features from point sequences and stroke distance fields, fusing them into a unified latent representation. This enriched embedding space facilitates smooth feature transitions during diffusion, enabling the generation of sketches with natural stroke fluidity, structural coherence, and fine-grained detail. Extensive qualitative and quantitative evaluations demonstrate the superiority of StrokeFusion in producing diverse, high-quality sketches across varying levels of complexity.
Acknowledgments
We thank the anonymous reviewers for their constructive feedback. This work was partially supported by grants from NSFC (62472287, U21B2023), Guangdong Basic and Applied Basic Research Foundation (2023A1515011297, 2023B1515120026), Shenzhen Natural Science Foundation (JCYJ20250604181519025), ICFCRT (W2441020), and the Scientific Development Funds from Shenzhen University.
@@ Acknowledgments
Figures 10 and 11 present additional qualitative results on both QuickDraw and several more complex sketch datasets. These visualizations further demonstrate the robustness and generalizability of our model across diverse domains.